OpenStudy (anonymous):

Simplify: x^2 - 9/ x^2 + 5x + 6 times 1/ x + 3

OpenStudy (btaylor):

factor them out, and cancel common factors \[\frac{(x+3)(x-3)}{(x+2)(x+3)(x+3)} \rightarrow \frac{x-3}{(x+2)(x+3)} \rightarrow \frac{x-3}{x^2+5x+6}\]